Frances is a family practice doctor who operates as a sole proprietor. She bought an X-ray...

  1. Frances is a family practice doctor who operates as a sole proprietor. She bought an X-ray machine 2 years ago for $50,000. She used the straight-line depreciation over the 5-year life of the machine. Today, Frances sold the machine for $58,000. In the current year, her marginal tax rate is 32%. How will Frances pay taxes on the gain from selling the X-ray machine?
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(50,000/5)=$10,000/year

Hence book value as on date of sale=Cost-Accumulated Depreciation

=50,000-(10,000*2)=$30,000

Hence gain on sale=(58,000-30,000)=$28,000

Hence taxes on gain from selling the X-ray machine=(28,000*32%)

=$8960